Candance has scores of 95, 92, 93, and 84 on her first 4 exams of the semester. what score must she obtain on the firth exam to have an average of a 90 or better for the 5 exams?
Answer by mananth(16949)

scores of 95, 92, 93, and 84 on her first 4 exams
TOTAL OF SCORES = 364
...
Average of should be 90
so total of 5 scores = 5*90= 450
..
Candance should score 450-364= 86
She should score 86 or more than 86 to get an average of 90 minimum